A building has a ramp to its front doors to accomodate the handicapped. If the distance from the building to the end of the ramp is 15 feet and the height from the ground to the fron doors is 6 feet, how long is the ramp?
A). 16.6 ft
B). 16.4 ft
c). 16.7 ft
D). 16.5 ft
E). 16.2 ft

Answer by sourbubble(2)

Your answer is d
x^2 = a^2 + b^2 Pythagoras
x^2 = 15^2 + 6^2
x^2 = 225 + 36
x^2 = 261
x = square root of 261
x = 16.155
therefore rounding the number gives you 6,2